a mam walks 15 blocks at a rate of 2 miles per hour . if there were 20 blocks in a mile, how long will it take him to walk
Answer by nycsharkman(136)

A mam walks 15 blocks at a rate of 2 miles per hour. If there were 20 blocks in a mile, how long will it take him to walk?
Use distance = rate (time) to find t.
15 = 2(t)
15/2 = t
7.5 = t
Set up a proportion:
7.5/t = 15/20
15x = 7.5 times 20
15x = 150
x = 150/15
x = 10 minutes